OpenStudy (anonymous):

What is the percent increase of 2,000 to 8,000. Show your work.

OpenStudy (anonymous):

Increase (8000 - 2000) 6000 Increase % = --------- * 100% = -------------- * 100% = ----- * 100 % original 2000 2000 Increase % = 3 * 100 % = 300%